# A Concept-Centric Approach to Multi-Modality Learning

### v1

@@ Abstract
In an effort to create a more efficient AI system, we introduce a new multi-modality learning framework that leverages a modality-agnostic concept space possessing abstract knowledge and a set of modality-specific projection models tailored to process distinct modality inputs and map them onto the concept space. Decoupled from specific modalities and their associated projection models, the concept space focuses on learning abstract knowledge that is universally applicable across modalities. Subsequently, the knowledge embedded into the concept space streamlines the learning processes of modality-specific projection models. We evaluate our framework on two popular tasks: Image-Text Matching and Visual Question Answering. Our framework achieves performance on par with benchmark models while demonstrating more efficient learning curves.
@@ 1 Introduction
Humans are capable of learning knowledge at a remarkable speed even during younger ages, which is in drastic contrast to most learning frameworks that require substantial resources to achieve human-like intelligence on specific tasks. Moreover, despite the exciting advancements from Large Language Models with multi-modality adaptations, there is hot debate over whether these models have achieved general intelligence or if they merely function via lossy compression of training corpora. We believe a concept-centric approach to multi-modality learning could be the key to not only bridging the efficiency gap but also marching towards a more natural learning process that mimics human learning.
At the center of our framework is a concept space that carries universal knowledge applicable to diverse modalities. Recent inspiring works on Concept Learning often focus on linking concepts to specific neurons (Liu et al., 2023b) and encoded embedding vectors (Kalibhat et al., 2023; Wang et al., 2023) of a model or injecting specific concepts as neurons into a model’s structure (Sheth & Kahou, 2023; Koh et al., 2020). Compared to these works, our proposed framework takes a systematic approach by organizing modality-agnostic abstract concepts in an interpretable knowledge space and establishing connections to different modalities by projecting modality-specific inputs onto the same space.
While it is common in multi-modality learning to create a shared representation space for multiple modalities (Radford et al., 2021; Li et al., 2022; Ramesh et al., 2022) or even utilize projections to align features from different modalities (Liu et al., 2023a), our shared concept space differentiates itself by possessing abstract knowledge which facilitates efficient learning and effortless incorporation of new modalities into the framework, as demonstrated in our experiments. We believe the proposed framework is a step closer to matching the capabilities of human learning, where we excel in creating a cohesive comprehension of concepts and seamlessly connecting multiple modalities, such as vision and language, to the learned knowledge.
@@
Specifically, as outlined in Fig. 1, the proposed multi-modality learning framework features an abstract concept space and a set of modality-specific projection models. The modality-agnostic concept space, inspired by prior works on structured embedding space (Vilnis et al., 2018; Li et al., 2018), optimally reflects real-world relations between concepts via entailment probabilities (Fig. 1 left). Probing into this concept space can also be achieved through simple queries of interested concept pairs, bringing interpretability into the learned knowledge.
Complementing the concept space, modality-specific projection models process and project distinct modality inputs onto the domain where the concept space is in (Fig. 1 middle). We call this domain the knowledge space as it hosts and bridges abstract knowledge embedded onto the concept space as well as specific knowledge extracted from material inputs. Decoupling the concept space from the projection models streamlines learning by unifying knowledge in an embedding space shared across modalities. The only restriction on the projection models is to produce consistent outputs that reside in the knowledge space, allowing them to be customized for their diverse modality inputs. This flexibility facilitates the seamless integration of different modalities, whose projection processes remain independent of each other. Naturally, the modular design of our framework extends its support to various downstream tasks, with inference processes conducted within the knowledge space (Fig. 1 right).
Contribution. Our contributions are three-fold. First, we propose a novel approach to multi-modality learning that centers around a concept space embedded with universally applicable knowledge. To our knowledge, this idea of a concept-focused learning scheme has rarely been explored in the field of multi-modality learning (Sec. 3). Second, we offer a clear motivation and justification for the proposed framework. Leveraging knowledge learned from the concept space, our framework demonstrates more efficient learning curves compared to traditional methods (Figure 2). The effectiveness of the concept space is further validated through an ablation study (Sec. 5). Third, we evaluate our framework’s performance on two downstream tasks. We show that the proposed framework, with a modest pretraining footprint, achieves comparable performance to benchmarks out-of-the-box without fine-tuning. (Sec. 4).
@@ 2 Related Work
Multi-Modality Learning. Vision and language modalities remain at the forefront of multi-modality learning research, with some works exploring alternative modalities like audio (Akbari et al., 2021; Shi et al., 2022). Within the vision-language area, CLIP by Radford et al. (2021) employs two modality-specific encoders to learn a joint representation through image-text matching. A subsequent work (Ramesh et al., 2022) introduces a text-to-image generation framework, using a text encoder and an image decoder for high-quality image generation from textual descriptions. Transformer-based architectures (Vaswani et al., 2017) have been widely explored for cross-modality information exchange and learning (Singh et al., 2022a; Bao et al., 2022; Kim et al., 2021a).
@@
Beyond combining and relating modalities, research has delved into diverse areas such as multi-modality few-shot learning (Alayrac et al., 2022; Li et al., 2021) and visual-textual pattern mining (He & Peng, 2020). Some studies propose generalized learning frameworks applicable across various modalities (Jaegle et al., 2021; Baevski et al., 2022a, b). While these frameworks showcase strong capabilities in tasks like text-to-image generation and visual language few-shot learning, our work addresses a distinct and important issue: creating a universally applicable concept space with abstract knowledge reflecting real-world observations. Baevski et al. present a versatile representation learning framework (2022b), yet it isolates modalities, impeding cross-modality interactions. In contrast, our proposed method directly combines modalities by projecting modality-specific inputs onto a unified concept space, eliminating the information barrier between them.
@@
Concept Learning. Early approaches to Concept Learning utilize Boolean logic for defining concepts based on relationships with other concepts (Angluin, 1988) and attributes (Mitchell, 1997). Lake et al. propose a Bayesian Program Learning framework, representing concepts as probabilistic programs (2015). Nowadays, a prevalent method involves placing concepts within a structured embedding space. Marconato et al. offer a clear interpretability definition for learned concepts in an embedding space (2022). Concept learning frameworks from Mao et al. (2019) and Li et al. (2020b) place similar concepts and their corresponding visual representations close to each other. Methods from Vilnis et al. (2018) and Mei et al. (2022) emphasize entailment relationships between concepts in learned concept spaces.
@@
In a departure from structured concept spaces, Liu et al. propose identifying "concept neurons" responsible for learning specific concepts in a deep net (2023b). While we acknowledge that some motivating works adopt a similar concept embedding strategy, our approach stands out for several reasons. The primary distinction lies in our concept space, which reflects real-world relations by providing meaningful numerical entailment probabilities that mirror those indicated by actual concepts. Furthermore, no barrier in our concept space prevents concepts belonging to different groups, such as red in color and cube in shape, from interacting with each other. Moreover, instead of being fitted to a specific modality, our concept space is designed to be abstract and modality-agnostic, which allows interactions between inputs from different modalities.
@@ 3 Method
Our proposed multi-modality learning framework consists of a modality-agnostic concept embedding space that models underlying relationships between concepts via entailment probabilities and a set of modality-specific projection models that extract representation from single-modality inputs and project them onto the domain where the concept space is in, i.e., the knowledge space.
Learning abstract knowledge in the concept space ensures generality, which makes its domain a good landing place for extracted representations from different modalities. Decoupled from the concept space and each other, modality-specific projection models can be tailored for adaptation to their unique inputs, while modality-specific knowledge stays connected after the projection.
@@ 3.1 Learning Concept Space
Davis et al. describe a knowledge representation as a surrogate that both carries the thing that exists in the real world and serves as a medium for pragmatically efficient computation (1993). Building upon their definition of a knowledge representation, we adopt an embedding space proposed by Li et al. (2018) to organize learned representations of abstract concepts. Like mental entities of specific knowledge in our brains, where we can relate concepts to each other, abstract entities in this concept space should be capable of interacting with each other, allowing reasoning inferences. In the proposed framework, we focus on entailment relations between concepts depicted by entailment probabilities to allow interactions between concepts. Contrary to latent spaces or learned ML model parameters, probing into the learned knowledge of this concept space can be easily achieved by querying the entailment probabilities of concept pairs of interest. Furthermore, our experiments demonstrate the efficiency of learning and referencing this concept space, facilitated by its compact parameter size, which qualifies it as a medium for pragmatically efficient computation.
@@
Defining Concept Space. We first define a knowledge space $\mathcal{K}\subset\mathbb{R}^{d}$ as a $d$-dimensional embedding space. Let $\mathcal{Y}$ be a set for modality-agnostic concepts. Each concept $y\in\mathcal{Y}$ is represented in $\mathcal{K}$ by a box embedding (the surrogate), defined by a pair of vectors $\Omega_{y}=(\omega_{min,y},\omega_{max,y})$, where $\omega_{min,y},\omega_{max,y}\in\mathcal{K}$ correspond to the minimum and maximum boundaries of the box in $\mathcal{K}$. We use $\mathcal{C}=\{\Omega_{y}|y\ \in\mathcal{Y}\}\subset\mathcal{K}$ to denote a set of box embeddings for every concepts in $\mathcal{Y}$ and we call $\mathcal{C}$ the concept space whose parameters are optimized to reflect real-world knowledge.
@@
A smoothing function $m_{\text{soft}}^{i}(\omega)=\frac{\text{softplus}(\omega^{i})}{\text{softplus}(G_{max}^{i}-G_{min}^{i})}$ is introduced on each dimension $i$ of $\mathcal{K}$ so a joint probability between two disjoint concepts can still be obtained. $G_{max}^{i},G_{min}^{i}$ terms are the global maximum and minimum values at the $i$ dimension among all $\Omega_{y}$s in $\mathcal{C}$. More details of $m_{\text{soft}}^{i}$ can be found in Appendix A.1. The probability of a single concept $y$ is calculated as $P(y)=P(\Omega_{y})=\prod_{i=1}^{d}m_{\text{soft}}^{i}(\omega_{max,y}-\omega_{min,y})$. The joint probability between two concepts $y_{1}$ and $y_{2}$ is calculated as $P(y_{1}\cap y_{2})=P(\Omega_{y_{1}}\cap\Omega_{y_{2}})=\prod_{i=1}^{d}m_{\text{soft}}^{i}(\text{min}(\omega_{max,y_{1}},\omega_{max,y_{2}})-\text{max}(\omega_{min,y_{1}},\omega_{min,y_{2}}))$.
@@
Modality-agnostic abstract knowledge can be extracted from $\mathcal{D}_{*}$ by examining entailment probabilities between concepts indicated by $\{\bm{y}_{i}\}_{i=1}^{N}$. Specifically, the ground-truth probability of a single concept and the entailment probability of a concept pair $(y_{1},y_{2})$ are calculated by $P(y)=\frac{\text{count}(y)}{\sum_{y^{\prime}\in\mathcal{Y}}\text{count}(y^{\prime})}$ and $P(y_{1}|y_{2})=\frac{\text{count}((y_{1}\cap y_{2}))}{\text{count}(y_{2})}$ as they appear in $\mathcal{D}_{*}$.
To drive the concept space to reflect real-world relationship between concepts via entailment probabilities, the objective for pretraining $\mathcal{C}$ is naturally defined as minimizing the Kullback–Leibler divergence between predicted probabilities obtained from $\mathcal{C}$ and true probabilities observed in $\mathcal{D_{*}}$. In addition to true concepts in $\bm{y}_{i}$ for each data point, a set of negative concepts is sampled and added to $\bm{y}_{i}$. A well-organized concept space should also reflect these negative concepts’ true entailment probabilities with the original concepts. Details of this negative sampling procedure vary by specific datasets and further information is provided in Sec. 4. For each sample, we calculate an entailment probability $Q(y_{1}|y_{2})$ indicated by the concept space for every possible combination of concept pairs $(y_{1},y_{2})$ in $\bm{y}_{i}$ and compare them to the true entailment probabilities $P(y_{1}|y_{2})$. We refer readers to Appendix A for further details regarding the concept space.

### v2

@@ Abstract
Humans possess a remarkable ability to acquire knowledge efficiently and apply it across diverse modalities through a coherent and shared understanding of the world. Inspired by this cognitive capability, we introduce a concept-centric multi-modality learning framework built around a modality-agnostic concept space that captures structured, abstract knowledge, alongside a set of modality-specific projection models that map raw inputs onto this shared space. The concept space is decoupled from any specific modality and serves as a repository of universally applicable knowledge. Once learned, the knowledge embedded in the concept space enables more efficient adaptation to new modalities, as projection models can align with existing conceptual representations rather than learning from scratch. This efficiency is empirically validated in our experiments, where the proposed framework exhibits faster convergence compared to baseline models. In addition, the framework’s modular design supports seamless integration of new modalities, since projection models are trained independently yet produce unified outputs within the shared concept space.
We evaluate the framework on two representative downstream tasks. While the focus is not on task-specific optimization, the framework attains comparable results with a smaller training footprint, no task-specific fine-tuning, and inference performed entirely within a shared space of learned concepts that offers interpretability. These findings point toward a promising direction for developing learning systems that operate in a manner more consistent with human cognitive processes.
@@ 1 Introduction
Humans are capable of acquiring knowledge at remarkable speed even from a young age, which stands in stark contrast to most learning frameworks that require substantial resources to achieve human-like intelligence on specific tasks. Moreover, human cognition is grounded in a shared and coherent understanding of the world that spans across different modalities. For instance, when learning a new language, we do not build an entirely separate system of knowledge for it. Instead, we intuitively connect new linguistic elements to our existing understanding of the world, or in other words, to our common sense. We believe a concept-centric approach to multi-modality learning could be key to not only bridging the efficiency gap, but also enabling learning frameworks that are more structured, reusable, and aligned with how humans organize knowledge across modalities.
At the center of our framework [^1] Code is available at https://github.com/Yuchong-Geng/concept-centric-multimodal-learning. is a concept space that carries universal knowledge applicable to diverse modalities, resembling the common sense embedded in the human mind. Recent inspiring works on Concept Learning often focus on linking concepts to specific neurons (Liu et al., 2023b) and encoded embedding vectors (Kalibhat et al., 2023; Wang et al., 2023b) of a model or injecting specific concepts as neurons into a model’s structure (Sheth and Kahou, 2023; Koh et al., 2020). Compared to these works, our proposed framework takes a systematic approach by organizing modality-agnostic abstract concepts in an interpretable knowledge space and establishing connections to different modalities by projecting modality-specific inputs onto the same space.
While it is common in multi-modality learning to create a shared representation space for multiple modalities (Radford et al., 2021; Li et al., 2022; Ramesh et al., 2022) or utilize projections to align features from different modalities (Liu et al., 2023a; Hsiung et al., 2022), such approaches primarily aim to co-locate modality-specific features in a common embedding space. In contrast, our shared concept space serves as an explicit intermediate layer of abstract concepts that can be reused and composed across modalities, enabling more efficient learning and seamless incorporation of new modalities, as demonstrated in our experiments. We believe the proposed framework takes a step toward matching key aspects of human learning, where we build cohesive concept-level understanding and connect multiple modalities (e.g., vision and language) to shared knowledge.
@@
Specifically, as outlined in Fig. 1, the proposed multi-modality learning framework features an abstract concept space and a set of modality-specific projection models. The modality-agnostic concept space, inspired by prior works on structured embedding spaces (Vilnis et al., 2018; Li et al., 2018), optimally reflects real-world relations between concepts via entailment probabilities (Fig. 1, left). Probing this concept space can also be achieved through simple queries of concept pairs of interest, providing interpretability into the learned knowledge.
Complementing the concept space, modality-specific projection models process inputs from different modalities and map them into a shared domain, which we refer to as the knowledge space (Fig. 1, middle). This knowledge space hosts both the abstract knowledge encoded in the concept space and the specific information extracted from individual inputs. By decoupling the projection models from the concept space, the framework enables efficient and modular learning. Each projection model is only required to produce consistent outputs within the knowledge space, allowing flexibility in architecture and optimization for different modalities. Although the projection models operate independently, their outputs are unified in the knowledge space, where they can interact with each other and with the learned concept representations, resulting in a structure that supports probabilistic reasoning and cross-modality interactions.
The proposed design, characterized by a shared concept space with universally applicable knowledge and flexible projection mechanisms, naturally facilitates the reuse of learned knowledge across diverse modalities and task domains. Such a design enhances the generalizability of our framework and enables straightforward adaptation to various downstream tasks, with all inference processes conducted within the knowledge space (Fig. 1, right).
Contribution. Our contributions are three-fold. First, we propose a novel approach to multi-modality learning that centers around a concept space embedded with universally applicable knowledge. To our knowledge, this idea of a concept-focused learning scheme has been underexplored in the field of multi-modality learning (Sec. 3). Second, we offer a clear motivation and justification for the proposed framework. Leveraging knowledge learned from the concept space, our framework demonstrates more efficient learning curves compared to traditional methods (Fig. 2). Additionally, although our projection models are trained independently, they still exhibit strong cross-modality alignment without joint training because they rely on the same concept space that encodes shared abstract knowledge. (Sec. 4.2) This design also makes the framework naturally scalable for incorporating new and diverse modalities (Sec. 4.3). Third, we evaluate the performance of our framework on two downstream tasks. We show that the proposed framework, with a modest pretraining footprint, achieves comparable performance to benchmarks out-of-the-box without fine-tuning, while conducting all inference within a shared knowledge space containing interpretable concept representations (Sec. 4.4).
@@ 2 Related Work
Multi-Modality Learning. Vision and language modalities remain at the forefront of multi-modality learning research, with some works exploring alternative modalities like audio (Akbari et al., 2021; Shi et al., 2022) and biomedical data (Masood et al., 2025). Within the vision-language area, CLIP by Radford et al. (2021) employs two modality-specific encoders to learn a joint representation through image-text matching. Subsequent work by Ramesh et al. (2022) introduces a text-to-image generation framework, using a text encoder and an image decoder to generate high-quality images from textual descriptions. Transformer-based architectures (Vaswani et al., 2017) have been widely explored for cross-modality information exchange and learning (Singh et al., 2022a; Bao et al., 2022; Kim et al., 2021a; Wang et al., 2023a; Lu et al., 2023).
@@
Beyond combining and relating modalities, research has delved into diverse areas such as multi-modality few-shot learning (Alayrac et al., 2022; Li et al., 2021) and visual-textual pattern mining (He and Peng, 2020). Some studies propose generalized learning frameworks applicable across various modalities (Jaegle et al., 2021; Baevski et al., 2022a; b). While these frameworks showcase strong capabilities in tasks like text-to-image generation and visual-language few-shot learning, our work addresses a distinct and important issue: creating a universally applicable concept space with abstract knowledge reflecting real-world observations. Baevski et al. (2022b) present a versatile representation learning framework, yet it isolates modalities, impeding cross-modality interactions. In contrast, our proposed method directly combines modalities by projecting modality-specific inputs onto a unified concept space, eliminating the information barrier between them.
@@
Concept Learning. Early approaches to Concept Learning utilized Boolean logic for defining concepts based on relationships with other concepts (Angluin, 1988) and their associated attributes (Mitchell, 1997). Lake et al. (2015) propose a Bayesian Program Learning framework that represents concepts as probabilistic programs. Nowadays, a prevalent method involves placing concepts within a structured embedding space. Concept learning frameworks such as those proposed by Mao et al. (2019) and Li et al. (2020b) construct embedding spaces that align concept representations with corresponding visual feature vectors. Lee et al. (2024) propose a framework that learns concept embeddings via distillation from pre-trained vision-language models. Methods from Vilnis et al. (2018) and Mei et al. (2022) emphasize entailment relationships between concepts in learned embedding spaces, while the work from Sinha et al. (2024) captures hierarchical information.
In a departure from structured concept embedding spaces, the Concept Bottleneck Model (CBM) (Koh et al., 2020) has become a popular framework that represents concepts as intermediate neural network outputs. CBM first predicts a set of pre-defined concepts aligned with human annotations and then produces a classification decision based on those concept predictions. Extending the idea of CBM, Dominici et al. (2023) propose a framework that learns a multimodal concept space jointly across modalities, using CBM-style supervision to inject interpretability into the shared concept space. Liu et al. (2023b) propose a method for identifying a small subset of model parameters responsible for generating specific concepts in a diffusion model. Kong et al. (2024) propose a theoretical view of concept learning as an identification problem of a discrete latent hierarchical model.
@@
While we acknowledge that some motivating works adopt a similar strategy involving a concept embedding space, our approach stands out for several reasons. The primary distinction lies in the organization of our concept space, which reflects real-world knowledge by providing meaningful numerical entailment probabilities that mirror relationships among actual concepts. Furthermore, no barrier in our concept space prevents concepts belonging to different groups, such as red in color and cube in shape, from interacting with each other. More importantly, instead of being fitted to a specific modality, our concept space is designed to be abstract and modality-agnostic, thus allowing interactions between inputs from different modalities.
@@ 3 Method
Our proposed multi-modality learning framework consists of a modality-agnostic concept embedding space that models underlying relationships between concepts via entailment probabilities and a set of modality-specific projection models that extract representation from single-modality inputs and project them onto the domain where the concept space resides, i.e., the knowledge space.
Learning abstract knowledge in the concept space ensures generality, which makes its domain a good landing place for extracted representations from different modalities. Decoupled from the concept space and each other, modality-specific projection models can be tailored for adaptation to their unique inputs, while modality-specific knowledge remains connected after the projection.
@@ 3.1 Learning Concept Space
Davis et al. (1993) describe a knowledge representation as a surrogate that both carries the thing existing in the real world and serves as a medium for pragmatically efficient computation. Building upon their definition of knowledge representation, we adopt a probabilistic box embedding space (Li et al., 2018) to organize the learned representations of abstract concepts. In this formulation, each concept is represented as an axis-aligned hyperrectangle (box) in a high-dimensional space, and geometric relations between boxes correspond directly to semantic relations. Intuitively, the strength of the relationship between two concepts is reflected in the amount of overlap between their boxes: a larger intersection indicates stronger semantic inclusion or entailment, while boxes with smaller overlap correspond to weaker or less related concepts.
@@
Like mental entities of specific knowledge in our brains, where we can relate concepts to each other, abstract entities in this concept space should be capable of interacting with each other, allowing reasoning inferences. With probabilistic box embeddings, these interactions take a geometric form: relations between concepts are expressed through the overlap structure of their boxes, which determines their entailment probabilities. In the proposed framework, we focus on entailment relations between concepts depicted by these entailment probabilities to allow interactions between concepts. Contrary to latent spaces or learned ML model parameters, probing into the learned knowledge of this concept space can be easily achieved by querying the entailment probabilities of concept pairs of interest. Furthermore, our experiments demonstrate the efficiency of learning and referencing this concept space, facilitated by its compact parameter size, which qualifies it as a medium for pragmatically efficient computation.
@@
Defining Concept Space. We first define a knowledge space $\mathcal{K}\subset\mathbb{R}^{d}$ as a $d$-dimensional embedding space. Let $\mathcal{Y}$ be a set for modality-agnostic concepts. Each concept $y\in\mathcal{Y}$ is represented in $\mathcal{K}$ by a box embedding (the surrogate), defined by a pair of vectors $\Omega_{y}=(\omega_{\text{min},y},\omega_{\text{max},y})$, where $\omega_{\text{min},y},\omega_{\text{max},y}\in\mathcal{K}$ correspond to the minimum and maximum boundaries of the box in $\mathcal{K}$. We use $\mathcal{C}=\{\Omega_{y}\mid y\ \in\mathcal{Y}\}\subset\mathcal{K}$ to denote a set of box embeddings for every concepts in $\mathcal{Y}$ and we call $\mathcal{C}$ the concept space whose parameters are optimized to reflect real-world knowledge.
@@
A smoothing function $m_{\text{soft}}^{i}(\omega)=\frac{\text{softplus}(\omega^{i})}{\text{softplus}(G_{\text{max}}^{i}-G_{\text{min}}^{i})}$ is introduced on each dimension $i$ of $\mathcal{K}$ so a joint probability between two disjoint concepts can still be obtained. $G_{\text{max}}^{i},G_{\text{min}}^{i}$ terms are the global maximum and minimum values at the $i$ dimension among all $\Omega_{y}$s in $\mathcal{C}$. More details of $m_{\text{soft}}^{i}$ can be found in Appendix A.1. The probability of a single concept $y$ is calculated as $P(y)=P(\Omega_{y})=\prod_{i=1}^{d}m_{\text{soft}}^{i}(\omega_{\text{max},y}-\omega_{\text{min},y})$. The joint probability between two concepts $y_{1}$ and $y_{2}$ is calculated as
$P(y_{1}\cap y_{2})=P(\Omega_{y_{1}}\cap\Omega_{y_{2}})=\prod_{i=1}^{d}m_{\text{soft}}^{i}(\text{min}(\omega_{\text{max},y_{1}},\omega_{\text{max},y_{2}})-\text{max}(\omega_{\text{min},y_{1}},\omega_{\text{min},y_{2}}))$
@@
Modality-agnostic abstract knowledge can be extracted from $\mathcal{D}_{*}$ by examining entailment probabilities between concepts indicated by $\{\bm{y}_{i}\}_{i=1}^{N}$. Specifically, the ground-truth probability of a single concept and the entailment probability of a concept pair $(y_{1},y_{2})$ are calculated by $P(y)=\frac{\text{count}(y)}{\sum_{y^{\prime}\in\mathcal{Y}}\text{count}(y^{\prime})}$ and $P(y_{1}\mid y_{2})=\frac{\text{count}((y_{1}\cap y_{2}))}{\text{count}(y_{2})}$ as they appear in $\mathcal{D}_{*}$.
To drive the concept space to reflect real-world relationships between concepts via entailment probabilities, the objective for pretraining $\mathcal{C}$ is naturally defined as minimizing the KL divergence between predicted probabilities obtained from $\mathcal{C}$ and the true probabilities observed in $\mathcal{D_{*}}$. In addition to the true concepts in $\bm{y}_{i}$ for each data point, we sample a set of negative concepts and append them to $\bm{y}_{i}$. This negative sampling is necessary because the supervision available in $\mathcal{D_{*}}$ contains only positive co-occurring concept pairs; without explicit counterexamples, the model would have no evidence that unrelated concepts should exhibit low or zero entailment. Negative samples therefore provide the missing contrastive signal, ensuring that the concept space learns to represent both strong entailments (via overlapping boxes) and non-entailments (via small or negligible overlap), rather than collapsing into an overly permissive representation. The specific sampling protocol varies across datasets, and further details are provided in Sec. 4.
For each sample and its set of concept labels $(x_{i}^{*},\bm{y}_{i})\in\mathcal{D}_{*}$, we compute the entailment probability $Q(y_{1}\mid y_{2})$ given by the concept space for every possible combination of concept pairs $(y_{1},y_{2})$ and for sampled negative concept pairs in $\bm{y}_{i}$, and we compare these values with the corresponding true entailment probabilities $P(y_{1}\mid y_{2})$. We refer readers to Appendix A for more details about the concept space.